The risk-free rate of interest, rf, is 6 percent. The overall stock market has an expected return of 12 percent. Maro, Inc. has a beta of 1.2. What is the required return of Maro, Inc. stock?
1
Expert's answer
2020-05-07T18:17:37-0400
  1. According to CAPM, expected return of a particular stock, for example, Maro, Inc. can be calculated as:

"E(r) = rf+\beta*(E(rm)-rf)"

where "E(r)" -expected return of a stock, "rf" - risk-free rate, "\beta" - beta, "E(rm)" - expected return of the overall market.

2. By plugging in given values we get:

"E(r) = rf+\beta*(E(rm)-rf) =\\\\= 6+ 1.2*(12-6)=13.2"

Required return of Maro is 13.2%